A simple compass can be made by placing a small iron bar magnet on a cork floating in the water. If the cork has a mass of 12.2 g and a 20-g magnet is placed on it, what fraction of the cork will be submerged?


Round your answer to 3 decimal places


1
Expert's answer
2021-04-19T17:07:41-0400

The cork will float when the buoyancy force is equal to the force of gravity:


"\rho_\text{water} g(V·k)=g(m+M)\\\space\\
V=\frac{m}{\rho_\text{cork}}\\\space\\
\rho_\text{water} g\frac{m}{\rho_\text{cork}}k=g(m+M),\\\space\\
k=\frac{(m+M)\rho_\text{cork}}{m\rho_\text{water}}=0.633."